A girl weighing 30kg skips 0.3 meter height 20 times calculate the work done by her
1
Expert's answer
2013-01-15T08:37:03-0500
The work done by a constant force of magnitude F ona point that moves a distance d in the direction of the forceis the product
A = F*d
In our case:
F = m*g
d = 20*0.3m = 6 m
A = m*g*6 m = 30*10*6 =1800 J